OpenStudy (meilendurcer):

Which ratios form a proportion? A. 4/5, 20/25 B. 8/12, 18/24 C. 1/3, 7/24 D. 2/5, 6/16

OpenStudy (anonymous):

20/25=4/5 4/5=4/5...hnce they are in proportion

OpenStudy (meilendurcer):

ok thanks

OpenStudy (anonymous):

hve u undrstood the concpt of proportion?

OpenStudy (meilendurcer):

not really

OpenStudy (anonymous):

in case of proportion...the ratio of no. on both side is same...

OpenStudy (meilendurcer):

hm .. ok now i know